OpenStudy (curry):

FIND THE equation of the tangent at x = 2 x^4 + x^3 + 1

OpenStudy (curry):

i got 4x^3+ 3x^2

OpenStudy (curry):

is that right

OpenStudy (bahrom7893):

y=x^4+x^3+1 y' = 4x^3 + 3x^2 y'(2) = 4*2^3 + 3*2^2

OpenStudy (bahrom7893):

and now just do: y-y(2) = y'(2)(x-2)

OpenStudy (curry):

wait i dont get that last part

OpenStudy (bahrom7893):

y-y1=M(x-x1)

OpenStudy (bahrom7893):

x1 = 2 y1 = y(2) = 2^4+2^3+1

OpenStudy (curry):

why + 1

OpenStudy (bahrom7893):

y = x^4 + x^3 + 1

OpenStudy (bahrom7893):

y(2) = 2^4+2^3+1

OpenStudy (curry):

oh oh kk

OpenStudy (curry):

so is the right answer y=(4x^3+3x^2)(x)+1

OpenStudy (curry):

wait i mean

OpenStudy (bahrom7893):

No... y-y(2) = y'(2)(x-2) y - 2^4 - 2^3 - 1 = (4*2^3+3*2^2)(x-2

OpenStudy (curry):

y = (4x^3+3x^2)(x)-63

OpenStudy (curry):

is that right?

OpenStudy (bahrom7893):

no, what i posted is. Just simplify it: y - 2^4 - 2^3 - 1 = (4*2^3+3*2^2)(x-2)

OpenStudy (curry):

but y is it y-y' i dont understand the ewquation

OpenStudy (curry):

i can do it i want to know y the equation is that way

OpenStudy (bahrom7893):

it's y-y(2) = y'(2)(x-2)

OpenStudy (curry):

right so if i was to simplify that

OpenStudy (bahrom7893):

y(2) is y of 2 not two times y

OpenStudy (curry):

i get y -25 = 4x^3+3x^2(x)-88

OpenStudy (bahrom7893):

btw you leave y as y..

OpenStudy (curry):

wat do u mean

OpenStudy (curry):

y = 4x^3+3x^2(x)-63 ??

OpenStudy (bahrom7893):

y - 2^4 - 2^3 - 1 = (4*2^3+3*2^2)(x-2) Just simplify the numbers here...

OpenStudy (curry):

ye thats wat i did

OpenStudy (curry):

but should i isolate y?

OpenStudy (bahrom7893):

then where are you getting 3x^2 and 4x^3 from......

OpenStudy (curry):

cause thats wat y'(2) is

OpenStudy (curry):

so y-y(2)=y'(2)(x-2)

OpenStudy (bahrom7893):

y'(2) is 4*2^3 + 3*2^2..

OpenStudy (bahrom7893):

y - 2^4 - 2^3 - 1 = (4*2^3+3*2^2)(x-2) y - 25 = 44(x-2)

OpenStudy (curry):

oo so then when the question asks for a derivative at a specific pooint then I would have to use what we did here but if it just says find the derivative of f(x) then i would put y = 4x^3+3x^2(x)

OpenStudy (curry):

or no just y = 4x^3 + 3x^2

OpenStudy (curry):

right?

OpenStudy (bahrom7893):

Nooooo.... you're not following y(x) means u plug in whatever x is into y.. example: y = 2x+1 x = 2 y(x) = 2*2+1 = 5

OpenStudy (curry):

ye i get that im just saying if it says in general

OpenStudy (curry):

to just din the derivative of some function and it is in rational format then i would use quotient rule to find the deravtive

OpenStudy (curry):

find* the derivative

OpenStudy (bahrom7893):

In general too. If you want to find an equation of the tangent line at x1 you find y'(x1). Then y(x1). Then you write: y-y(x)=y'(x)(x-x1)

OpenStudy (bahrom7893):

and that's it

OpenStudy (curry):

right right i was talking about just a derivative not a tangent line

OpenStudy (bahrom7893):

yea